# Human-Like Robot Impedance Regulation Skill Learning from Human-Human Demonstrations

### v1

@@ V Conclusions
In this paper, we propose a novel learning from demonstration framework to acquire impedance regulation skills from human-human collaboration. By calibrating the human dynamic model, the endpoint stiffness can be estimated. The reference trajectories are then calculated based on the interaction model with these measured data. A TP-LQT method is proposed to represent human movements and generate a new reference trajectory for the cobot under new situation. Besides, the impedance regulation skills of each task are learned through the LSTM networks to construct the regulation modules. Finally, we apply a human-like cobot CURI to human-robot collaboration tasks that perform object transportation and Tai Chi pushing hands with a human partner. We propose a whole-body impedance controller for this unique cobot and put forward a variable impedance control method based on the human-human demonstration. The real-world platform we built successfully achieved the collaborative object transportation task and Tai Chi pushing hands tasks, which will be a foundation of future research in HRC. The overall framework we propose to let the cobots learn impedance regulation skills from human-human demonstration can be extended to other tasks with variable physical behaviors and repetitive actions. For instance, the collaborative carrying of objects in logistic scenarios and the assistive dressing of clothes in domestic scenarios are feasible applications to achieve.
In this study, we limited our analysis to only the arms and torso of our cobot during collaboration, which constrains the cobot workspace and functional capacity to some extent. Additionally, considerations of human safety and physical ergonomics were not addressed by the cobot. To enhance task execution, our future work will incorporate a capture system that includes feedback on human kinematics and dynamics. Moreover, the current setup for human-human demonstrations is complex, and the transferability of learned behaviors from one task to another remains underdeveloped. We plan to investigate novel data representations to extract richer and more relevant interaction information during collaborations, aiming to optimize the efficiency of demonstrations.

### v2

@@ V Conclusions
In this paper, we propose a novel HI-ImpRSL framework to acquire impedance regulation skills from human-human collaboration. By calibrating the human dynamic model, we estimate endpoint stiffness and then compute reference trajectories based on the interaction model with these measured data. A TP-LQT method is proposed to represent human movements and generate a new reference trajectory for the cobot under new situations. Besides, the impedance regulation skills of each task are learned through the LSTM networks to construct the regulation modules. Finally, we apply a human-like cobot CURI to HRC tasks to perform object transportation and Tai Chi pushing hands with a human partner. We propose a whole-body impedance controller for this unique cobot and put forward a variable impedance control method based on the human-human demonstrations. The real-world platform we built successfully achieved the collaborative object transportation task, Tai Chi pushing hands tasks, and collaborative sawing tasks, demonstrating effective adaptive impedance regulation. This study bridges human cognitive behaviors and robotic skill learning, advancing HRC and providing a foundation for future research in robotics.
Nevertheless, we limited our analysis to only the arms and torso of our cobot during collaboration, which constrains the cobot workspace and functional capacity to some extent. Additionally, considerations of human safety and physical ergonomics were not addressed by the cobot. To enhance task execution, our future work will incorporate a capture system that includes feedback on human kinematics and dynamics. Moreover, the current setup for human-human demonstrations is complex, and the transferability of learned behaviors from one task to another remains underdeveloped. Future research will focus on designing more robust, generalized data representations inspired by cognitive systems, allowing the framework to extract richer interaction patterns and improve its adaptability to diverse collaborative scenarios.